Consider a market for steel.

A) Suppose that the supply curve for steel (in tonnes) is given by P = 2 + 0.01Q.'

i. Calculate the price elasticity of supply if the price is equal to $300 per tonne.

ii. If the price goes up, what do you expect will happen to the elasticity of supply. Explain.

iii. Explain and give examples of two factors that would cause the elasticity of supply for steel to increase.
Answer #1

Supply curve for steel is P = 2 + 0.01Q. When P = 300, Q = 29800 tonnes

i. Price elasticity of supply = 1/slope of supply function x P/Q

= 1/0.01 x 300 / 29800

= 1.006 or approximately 1.

ii. If the price goes up, the quantity supplied will also increase but the same is reduced in percentage terms so that the elasticity would gradually fall. This is because along the supply curve, as quantity is increased (or as price increases), elasticity of supply decreases.

iii. Elasticity of supply for steel depends on the time period. In the short run quantity supplied cannot be increased so supply remains inelastic relative to long run when the factors of production are variable and production can be increased. Also, the technique of production can be complex, thereby reducing the ease with which production is carried out. This makes supply inelastic. If the production technique is altered and it is made simpler, the supply can be made elastic (elasticity increases).